# Multimodal Datasets with Controllable Mutual Information

### v1

@@ Abstract
We introduce a framework for generating highly multimodal datasets with explicitly calculable mutual information between modalities. This enables the construction of benchmark datasets that provide a novel testbed for systematic studies of mutual information estimators and multimodal self-supervised learning techniques. Our framework constructs realistic datasets with known mutual information using a flow-based generative model and a structured causal framework for generating correlated latent variables. [^1] Our code is publicly available at https://github.com/RKHashmani/MmMi-Datasets.
@@ 1 Introduction
Self-supervised learning (SSL) has become a core component of many state-of-the-art large-scale machine learning models [1]. Such models are also increasingly multimodal, i.e. designed to learn from varied input sources such as text, images, and audio [2, 3]. A prevailing intuition is that multimodal SSL is effective because different modalities provide complementary “views” of the same underlying concepts, enabling the learning process to exploit their shared information. The precise relationship between the mutual information (MI) between modalities and SSL performance, however, is not fully understood.
Contrastive SSL methods built using the InfoNCE loss function [4, 5, 6, 7] have a clear information-theoretic interpretation: for example, the learned similarity scores estimate the pointwise mutual information (PMI) between paired samples. By contrast, no analogous theoretical connection has been established for either highly multimodal settings (i.e. $N>2$ modalities) or for non-contrastive SSL methods such as multimodal masked modeling [8, 9, 10, 11], despite the fact that both of these directions are quickly gaining prominence in the field [12, 13]. A theoretically-grounded understanding of the fundamental relationship between inter-modality mutual information and SSL representations (and their corresponding performance on downstream tasks) will be increasingly critical as models continue to scale to larger numbers of input modalities. In particular, principled frameworks will be needed to evaluate how the distribution of shared information across modalities influences the quality of the learned embeddings.
Complicating matters further, MI is notoriously difficult to estimate from samples, particularly in high-dimensional, real-world datasets [14, 15]. A wide range of MI estimators have been proposed using techniques such as kernel estimation, $k$-nearest neighbor, and neural estimators [16, 17, 18, 19, 20, 21, 22, 23]. However, these estimators are typically only validated on synthetic datasets of simple distributions for which the MI is analytically tractable [24, 25, 15, 26, 27].
@@
In this work, we introduce a framework to generate realistic multimodal data with controllable mutual information. Figure 1 shows an overview of our data generation framework:
@@ 2 Background
When evaluated on specific values $x\sim p(X)$ and $y\sim p(Y)$, the pointwise MI (PMI) captures the probability of these two values occurring together compared with that same probability if they were fully independent. The PMI is formally expressed as:
@@
Flow-based generative models are designed to facilitate the direct transformation between probability distributions using invertible mappings applied to a simple base distribution such as a Gaussian. Each transformation is designed to be bijective with a tractable Jacobian determinant, enabling exact computation of both likelihoods and samples. Because they provide both efficient sampling and exact density evaluation, flow-based models are increasingly used not only in generative modeling but also in scientific applications requiring tractable likelihoods and explicit control over distributions. Flow-matching [28, 29] is a recent approach within the family of flow-based generative models that enables efficient training of Continuous Normalizing Flows (CNFs) [30] by directly regressing the velocity field that transports a base distribution to the data distribution instead of optimizing the exact maximum-likelihood objective.
@@ 3.1 Generalized Linear Causal Construction: Proto-latent to Latent Connections
We wish to create a large latent variable vector $\mathbf{z}=(z_{\theta},\mathbf{z}_{1},\dots,\mathbf{z}_{N_{z}})^{T}$ that is distributed according to a multivariate Gaussian with known covariance for which the mutual information is easy to compute. We achieve this by forming linear combinations of i.i.d. normally distributed proto-latents $\mathbf{u}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$:
$\mathbf{z}=\begin{pmatrix}z_{\theta}\,\\ \mathbf{z}_{1}\\ \vdots\\ \,\mathbf{z}_{N_{z}}\end{pmatrix}=\mathbf{A}\begin{pmatrix}\mathbf{\tilde{u}}\,\,\\ \mathbf{\hat{u}}_{1}\\ \vdots\\ \,\mathbf{\hat{u}}_{N_{u}}\end{pmatrix}$ (1)
@@
  • $\mathbf{A}$: a user-defined matrix specifying the structural equations in the causal model relating $\mathbf{u}$ and $\mathbf{z}$.
@@
$z_{\theta}$ $=\eta\tilde{u}_{1}\;$
[…]
$\mathbf{z}_{2}$ $=\tilde{\rho}_{12}\tilde{{u}}_{1}\,\mathbf{1}_{d}+\hat{\rho}_{22}\hat{\mathbf{u}}_{2}\;,$
@@ 3.2 Analytic Mutual Information of the Latent Variables
We can represent the covariance matrix in block form corrresponding to $z_{\theta}$, $\mathbf{z}_{1}$, and $\mathbf{z}_{2}$ as
$\Sigma=\begin{bmatrix}\Sigma_{\theta\theta}&\Sigma_{\theta 1}&\Sigma_{\theta 2}\\ \Sigma_{1\theta}&\Sigma_{11}&\Sigma_{12}\\ \Sigma_{2\theta}&\Sigma_{21}&\Sigma_{22}\\ \end{bmatrix}$ (4)
For any two blocks in $\Sigma$, we define the reduced block matrix:
$\Gamma_{ij}=\begin{bmatrix}\Sigma_{ii}&\Sigma_{ij}\\ \Sigma_{ji}&\Sigma_{jj}\end{bmatrix}$ (5)
@@
$I(\theta;Z_{1})$ $=\frac{1}{2}\ln\left(\frac{|\Sigma_{\theta\theta}||\Sigma_{11}|}{|\Gamma_{\theta 1}|}\right)$ (6)
$I(Z_{1};Z_{2})$ $=\frac{1}{2}\ln\left(\frac{|\Sigma_{11}||\Sigma_{22}|}{|\Gamma_{12}|}\right)\;,$ (7)
@@
While the covariance matrix and mutual information quantities in the preceding equations can be calculated numerically, we are also able to derive closed-form, analytical equations for various mutual information quantities (see Appendix A.1). One benefit of the closed form solutions is they reveal scaling in terms of the hyperparameters of the structural equations, the number of modalities, and the dimensonality of each modality. In the case of the causal model considered in Fig. 1(a) and Eq. 3.1, we find
@@ 3.3 Flow-based generative modeling preserves mutual information
This result can be seen as following from the data-processing inequality and is also the result of a direct computation of the mutual information after a change of variables, where the Jacobian factors that arise cancel exactly [see e.g., 31, 15, 26]. While many generative models satisfy this condition, e.g., discrete-time normalizing flows [32, 33], we use continuous-time normalizing flows based on flow matching  [29, 28, 34] in this work. We pretrain on CIFAR-10 [35] using image class as a proxy for modality (i.e., $f_{0}$ is trained on images of cars, $f_{1}$ is trained on images of frogs), but we emphasize that our framework is agnostic to $f_{i}$ parameterization and modality definition.
@@ 4 Examples of datasets resulting from our model
Figure 2: Eight examples of correlated pairs of images $(\mathbf{x}_{1},\mathbf{x}_{2})$ generated from our procedure representing two realistic modalities. In this case, both modalities are images, but corresponding to flows conditioned on different class labels (“automobile”, “frog”) from CIFAR-10 [35]. The dimensionality of the data in both cases is $d=32\times 32\times 3=3072$.
@@
Figure 3: Examples of causal structures with corresponding linear structural equations that induce specific mutual information.
4.1 Example 1: Estimating a black hole’s mass from two telescopes
Consider the hypothetical scenario where one wishes to estimate the mass of Sagittarius A*, the supermassive black hole in the center of our Milky Way Galaxy. To do this we might employ two instruments producing two data modalities. Let $\mathbf{x}_{1}$ represent data from the Event Horizon Telescope, a ground-based array consisting of a global network of radio telescopes. Let $\mathbf{x}_{2}$ represent data from the Hubble space telescope in orbit around the Earth. Let $\tilde{u}_{1}$ represent the unknown mass of the black hole and let $\tilde{u}_{2}$ represent some atmospheric variability that impacts how radio waves propagate in the atmosphere.
@@ 4.1 Example 1: Estimating a black hole’s mass from two telescopes
$z_{\theta}$ $=\eta_{1}\tilde{u}_{1}+\eta_{2}\tilde{u}_{2}\;$
$\mathbf{z}_{1}$ $=\tilde{\rho}_{11}\tilde{u}_{1}\mathbf{T}_{11}+\tilde{\rho}_{12}\tilde{u}_{2}\mathbf{T}_{12}+\hat{\rho}_{11}\mathbf{\hat{u}}_{1}\;$ (12)
$\mathbf{z}_{2}$ $=\tilde{\rho}_{21}\tilde{u}_{1}\mathbf{T}_{21}+\hat{\rho}_{22}\mathbf{\hat{u}}_{2}\;.$
The closed-form, analytical equations for various mutual information quantities corresponding to a similar causal model can be found in Appendix A.1. The $\mathbf{T}_{ki}$ are templates that have the same shape as the $\mathbf{z}_{i}$ and can encode some type of inhomogeneous (spatial) structure in the latents. For example, the templates $\mathbf{T}_{11}$ and $\mathbf{T}_{12}$ associated with $\tilde{u}_{1}$ (black hole mass) are designed to concentrate at the center of the galaxy and dissipates away from the center. Similarly, the template $\mathbf{T}_{21}$ associated with $\tilde{u}_{2}$ (atmospheric effect) are designed to be diffuse across the whole example. Not shown explicitly in the figure are the functions that generate the observed data from the latents: $\theta=z_{\theta}$, $\mathbf{x_{1}}=f_{1}(\mathbf{z}_{1})$, and $\mathbf{x_{2}}=f_{2}(\mathbf{z}_{2})$.
@@ 4.2 Example 2: A scalable model for massively multimodal data
We consider the causal model illustrated in Fig. 3(b). This causal model corresponds to the structural equations
$z_{\theta}=\eta\tilde{u}_{1}\;\hskip 28.45274pt\mathbf{z}_{i}=\beta^{-i}\tilde{\rho}\,\tilde{u}_{1}\mathbf{1}+\sum_{j=1}^{N_{u}}\alpha^{-|i-j|}\hat{\rho}\,\mathbf{\hat{u}}_{j}\;.$ (13)
@@
(a) Mutual information between image modalities $X_{1}$ and $X_{i}$, for the $i$-th modality.
(b) Mutual information between $\theta$ and image
modality $X_{i}$, for the $i$-th modality.
Figure 4: Information between image modalities $X_{1}$ and $X_{i}$ decreases as the distance between $X_{1}$ and $X_{i}$ increases. This, as well as the information between image modality $X_{i}$ and the parameter $\theta$, decreases as the total number of modalities increases.
@@ 4.3 Example 3: A model for ablation studies for multimodal SSL
As discussed in Section 3.4, templates can control how the information is distributed among the components of each latent $\mathbf{z}_{i}$ while preserving the total mutual information between two modalities. This provides a mechanism for disentangling the effects of algorithmic (e.g. specific SSL objectives) from architectural choices (e.g. inductive biases in the model construction) by independently varying where the information is distributed in the data. The following structural equations incorporate modality-specific templates associated to a set of shared proto-latents representing common causes $\tilde{u}_{k}$ as well as a path for shared information from the proto-latents $\mathbf{\hat{u}}_{j}$ that are independent of the target latent $z_{\theta}$:

### v2

@@ Abstract
We introduce a framework for generating highly multimodal datasets with explicitly calculable mutual information (MI) between modalities. This enables the construction of benchmark datasets that provide a novel testbed for systematic studies of mutual information estimators and multimodal self-supervised learning (SSL) techniques. Our framework constructs realistic datasets with known MI using a flow-based generative model and a structured causal framework for generating correlated latent variables. We benchmark a suite of MI estimators on datasets with varying ground truth MI values and verify that regression performance improves as the MI increases between input modalities and the target value. Finally, we describe how our framework can be applied to contexts including multi-detector astrophysics and SSL studies in the highly multimodal regime. [^1] Our code is publicly available at https://github.com/RKHashmani/MmMi-Datasets.
@@ 1 Introduction
Self-supervised learning (SSL) has become a core component of many state-of-the-art large-scale machine learning models [balestriero2023cookbookselfsupervisedlearning]. Such models are also increasingly multimodal, i.e. designed to learn from varied input sources such as text, images, and audio [radford2021learning, zong2024selfsupervisedmultimodallearningsurvey]. A prevailing intuition is that multimodal SSL is effective because different modalities provide complementary “views” of the same underlying concepts, enabling the learning process to exploit their shared information. The precise relationship between the mutual information (MI) between modalities and SSL performance, however, is not fully understood.
Contrastive SSL methods built using the InfoNCE loss function [oord2018representation, chen2020simple, he2020momentum, caron2020unsupervised] have a clear information-theoretic interpretation: for example, the learned similarity scores estimate the pointwise mutual information (PMI) between paired samples. By contrast, no analogous theoretical connection has been established for either highly multimodal settings (i.e. $N>2$ modalities) or for non-contrastive SSL methods such as multimodal masked modeling [mizrahi20234m, he2022masked, wang2023image, huang2025learning], despite the fact that both of these directions are quickly gaining prominence in the field [li2024maskedmodelingselfsupervisedrepresentation, hondru2025maskedimagemodelingsurvey]. A theoretically-grounded understanding of the fundamental relationship between inter-modality mutual information and SSL representations (and their corresponding performance on downstream tasks) will be increasingly critical as models continue to scale to larger numbers of input modalities. In particular, principled frameworks will be needed to evaluate how the distribution of shared information across modalities influences the quality of the learned embeddings.
Complicating matters further, MI is notoriously difficult to estimate from samples, particularly in high-dimensional, real-world datasets [McAllester-2020-FormalLimitations, czyz2023beyond]. A wide range of MI estimators have been proposed using techniques such as kernel estimation, $k$-nearest neighbor, and neural estimators [pizer1987adaptive, kozachenko:87:firstnn, moon:95:kde, kraskov:04:ksg, belghazi:18:mine, song:20:understanding, butakov2024mutual, belghazi2021minemutualinformationneural]. However, these estimators are typically only validated on synthetic datasets of simple distributions for which the MI is analytically tractable [darbellay:99:adaptive-partitioning, suzuki:16:jic, czyz2023beyond, czyz2023properties, butakov2023information].
@@
In this work, we introduce a framework to generate realistic multimodal data with controllable mutual information. Figure 1 shows an overview of our data generation framework [^2] Datasets generated based on Figure 1 can be found at [raheem_hashmani_2026].:
@@ 2 Background
When evaluated on specific values $x\sim p(X)$ and $y\sim p(Y)$, the pointwise MI (PMI) captures the probability of $x$ and $y$ occurring together compared with that same probability if they were fully independent. The PMI is formally expressed as:
@@
Flow-based generative models are designed to facilitate the direct transformation between probability distributions using invertible mappings applied to a simple base distribution such as a Gaussian. Each transformation is designed to be bijective with a tractable Jacobian determinant, enabling exact computation of both likelihoods and samples. Because they provide both efficient sampling and exact density evaluation, flow-based models are increasingly used not only in generative modeling but also in scientific applications requiring tractable likelihoods and explicit control over distributions. Flow-matching [Albergo2022BuildingNF, lipman2024flow] is a recent approach within the family of flow-based generative models that enables efficient training of Continuous Normalizing Flows (CNFs) [chen2018continuoustimeflowsefficientinference] by directly regressing the velocity field that transports a base distribution to the data distribution instead of optimizing the exact maximum-likelihood objective.
@@ 3.1 Generalized Linear Causal Construction: Proto-latent to Latent Connections
We wish to create a large latent vector $\mathbf{z}$ that is distributed according to a multivariate Gaussian with known covariance for which the mutual information is easy to compute. We achieve this with linear combinations of i.i.d. normally distributed proto-latents $\mathbf{u}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$:
$\mathbf{z}=(z_{\theta},\mathbf{z}_{1},\dots,\mathbf{z}_{N_{z}})^{\top}=\mathbf{A}\,(\tilde{\mathbf{u}},\hat{\mathbf{u}}_{1},\cdots,\hat{\mathbf{u}}_{N_{u}})^{\top}$ (1)
@@
  • $\mathbf{A}$: a matrix specifying the structural equations in the causal model relating $\mathbf{u}$ and $\mathbf{z}$.
@@
$z_{\theta}$ $=\eta\tilde{u}_{1}\;;\hskip 30.00005pt\mathbf{z}_{1}=\tilde{\rho}_{11}\tilde{{u}}_{1}\,\mathbf{1}_{d}+\hat{\rho}_{11}\hat{\mathbf{u}}_{1}\;;\hskip 30.00005pt\mathbf{z}_{2}=\tilde{\rho}_{12}\tilde{{u}}_{1}\,\mathbf{1}_{d}+\hat{\rho}_{22}\hat{\mathbf{u}}_{2}\;,$ (2)
@@ 3.2 Analytic Mutual Information of the Latent Variables
We can represent the covariance matrix in block form corresponding to $z_{\theta}$, $\mathbf{z}_{1}$, and $\mathbf{z}_{2}$ as $\Sigma$ (and for any two blocks in $\Sigma$, we can define the reduced block matrix $\Gamma$) as:
$\Sigma=\begin{bmatrix}\Sigma_{\theta\theta}&\Sigma_{\theta 1}&\Sigma_{\theta 2}\\ \Sigma_{1\theta}&\Sigma_{11}&\Sigma_{12}\\ \Sigma_{2\theta}&\Sigma_{21}&\Sigma_{22}\\ \end{bmatrix}\qquad\Gamma_{ij}=\begin{bmatrix}\Sigma_{ii}&\Sigma_{ij}\\ \Sigma_{ji}&\Sigma_{jj}\end{bmatrix}$ (4)
@@
$I(\theta;Z_{1})$ $=\frac{1}{2}\ln\left(\frac{|\Sigma_{\theta\theta}||\Sigma_{11}|}{|\Gamma_{\theta 1}|}\right)\qquad I(Z_{1};Z_{2})=\frac{1}{2}\ln\left(\frac{|\Sigma_{11}||\Sigma_{22}|}{|\Gamma_{12}|}\right)\;,$ (5)
@@
While the covariance matrix and mutual information quantities in the preceding equations can be calculated numerically, we are also able to derive closed-form, analytical equations for various mutual information quantities (see Appendix B). One benefit of the closed form solutions is they reveal scaling in terms of the hyperparameters of the structural equations, the number of modalities, and the dimensonality of each modality. For the causal model considered in Fig. 1(a) and Eq. 2, we find:
@@ 3.3 Flow-based generative modeling preserves mutual information
This result can be seen as following from the data-processing inequality and is also the result of a direct computation of the mutual information after a change of variables, where the Jacobian factors that arise cancel exactly [see e.g., cover2006elements, czyz2023beyond, czyz2023properties]. While many generative models satisfy this condition, e.g., discrete-time normalizing flows [rezende2015variational, papamakarios2021normalizing], we use continuous-time normalizing flows based on flow matching  [lipman2024flow, Albergo2022BuildingNF, liu2022flow] in this work. We pretrain on CIFAR-10 [cifar10] using image class as a proxy for modality (i.e., $f_{0}$ is trained on images of cars, $f_{1}$ is trained on images of frogs), but we emphasize that our framework is agnostic to $f_{i}$ parameterization and modality definition.
@@ 4 Examples of datasets resulting from our model
Figure 2: Eight examples of correlated pairs of images $(\mathbf{x}_{1},\mathbf{x}_{2})$ generated from our procedure representing two realistic modalities. In this case, both modalities are images, but corresponding to flows conditioned on different class labels (“automobile”, “frog”) from CIFAR-10 [cifar10]. The dimensionality of the data in both cases is $d=32\times 32\times 3=3072$.
@@
4.1 Example 1: Empirically Demonstrating Example Use Cases
Benchmarking mutual information estimators.
We demonstrate the use of our framework to generate a benchmark dataset to explore the performance of a number of popular mutual information estimators. We use the causal model shown in Figure 1 to generate a set of ten datasets with mutual information $I(X_{1};X_{2})$ ranging from $0.0284$ to $1.39$, with each dataset consisting of 10,000 paired CIFAR-like images. We use an existing benchmark suite [lee2024benchmarksuiteevaluatingneural] to estimate the empirical mutual information and compare it to the ground-truth mutual information in Figure 4. We report the correlation and RMSE for each estimator in Table 1. We observe that the regressed mutual information from all the estimators follows the linearly increasing ground truth mutual information.
MI Estimator Correlation RMSE
DV [Donsker1983] 0.995 0.1094
JS [fGAN] 0.993 0.3049
InfoNCE [oord2018representation] 0.991 0.1981
MINE [belghazi:18:mine] 0.993 0.0983
NWJ [Nguyen_2010] 0.996 0.0851
SMILE-1 [SMILE] 0.998 0.0935
SMILE-5 [SMILE] 0.993 0.0974
SMILE-inf [SMILE] 0.993 0.0969
Table 1: Correlation and RMSE from linear regression of estimated vs. ground-truth mutual information, computed for a number of mutual information estimators.
Regressing the target variable $\theta$ from $X$.
We evaluate how well a model is able to regress the target variable $\theta$ from data $X_{1}$ as a function of the mutual information $I(\theta,X_{1})$. Intuitively, models with fixed capacity should predict $\theta$ more accurately from data that contain more information about $\theta$, i.e. larger $I(\theta;X_{1})$. We re-use the same ten datasets, for which $I(\theta;X_{1})$ ranges from $0.134$ to $1.73$. For each dataset, we regress $\theta$ from the images using a shallow convolutional network, choosing the best model after training for $500$ epochs. We show that the best achievable RMSE decreases with increasing mutual information between $X_{1}$ and $\theta$ (Figure 4) and show example distributions of prediction error $\theta-\hat{\theta}$ for two representative MI values in Appendix A.
Figure 3: MI estimators vs. ground-truth MI.
Figure 4: Regression of a target scalar $\theta$ monotonically improves with increasing $I(X_{1},\theta)$.
4.2 Example 2: Estimating a black hole’s mass from two telescopes
Consider the hypothetical scenario where one wishes to estimate the mass of Sagittarius A*, the supermassive black hole in the center of our Milky Way Galaxy. To do this we might employ two instruments producing two data modalities. Let $\mathbf{x}_{1}$ represent data from the Event Horizon Telescope, a ground-based array consisting of a global network of radio telescopes. Let $\mathbf{x}_{2}$ represent data from the Hubble space telescope in orbit around the Earth. Let $\tilde{u}_{1}$ represent the unknown mass of the black hole and let $\tilde{u}_{2}$ represent some atmospheric variability that impacts how radio waves propagate.
@@ 4.2 Example 2: Estimating a black hole’s mass from two telescopes
Figure 5: Example causal structures with corresponding structural equations that induce specific MI.
@@
$z_{\theta}$ $=\eta_{1}\tilde{u}_{1}+\eta_{2}\tilde{u}_{2}\;;\hskip 10.00002pt\mathbf{z}_{1}=\tilde{\rho}_{11}\tilde{u}_{1}\mathbf{T}_{11}+\tilde{\rho}_{12}\tilde{u}_{2}\mathbf{T}_{12}+\hat{\rho}_{11}\mathbf{\hat{u}}_{1}\;;\hskip 10.00002pt\mathbf{z}_{2}=\tilde{\rho}_{21}\tilde{u}_{1}\mathbf{T}_{21}+\hat{\rho}_{22}\mathbf{\hat{u}}_{2}\;.$ (10)
The closed-form, analytical equations for various MI quantities corresponding to a similar causal model can be found in Appendix B. The $\mathbf{T}_{ki}$ are templates that have the same shape as the $\mathbf{z}_{i}$ and can encode some type of inhomogeneous (spatial) structure in the latents. For example, the templates $\mathbf{T}_{11}$ and $\mathbf{T}_{12}$ associated with $\tilde{u}_{1}$ (black hole mass) are designed to concentrate at the center of the galaxy and dissipates away from the center. Similarly, the template $\mathbf{T}_{21}$ associated with $\tilde{u}_{2}$ (atmospheric effect) are designed to be diffuse across the whole example.
@@ 4.3 Example 3: A scalable model for massively multimodal data
The causal model illustrated in Fig. 5(b) corresponds to the structural equations
$z_{\theta}=\eta\tilde{u}_{1}\;;\hskip 28.45274pt\mathbf{z}_{i}=\beta^{-i}\tilde{\rho}\,\tilde{u}_{1}\mathbf{1}+\sum_{j=1}^{N_{u}}\alpha^{-|i-j|}\hat{\rho}\,\mathbf{\hat{u}}_{j}\;.$ (11)
@@
Figure 6: As expected, $I(X_{1},X_{i})$ (left) and $I(\theta,X_{i})$ (right) decrease with the index ii.
@@ 4.4 Example 4: A model for ablation studies for multimodal SSL
As discussed in Section 3.4, templates can control how the information is distributed among the components of each latent $\mathbf{z}_{i}$ while preserving the total mutual information between two modalities. This provides a mechanism for disentangling the effects of algorithmic (e.g. specific SSL objectives) from architectural choices (e.g. inductive biases in the model construction) by independently varying how the information is distributed in the data. The following equations incorporate modality-specific templates associated to a set of shared proto-latents representing common causes $\tilde{u}_{k}$ as well as a path for shared information from the proto-latents $\mathbf{\hat{u}}_{j}$ that are independent of the target latent $z_{\theta}$: